# SmartUT: Receive Beamforming for Spectral Coexistence of NGSO Satellite Systems

### v1

@@ Abstract
In this paper, we investigate downlink co-frequency interference (CFI) mitigation in non-geostationary satellites orbits (NGSOs) co-existing systems. Traditional mitigation techniques, such as Zero-forcing (ZF), produce a null towards the direction of arrivals (DOAs) of the interfering signals, but they suffer from high computational complexity due to matrix inversions and required knowledge of the channel state information (CSI). Furthermore, adaptive beamformers, such as sample matrix inversion (SMI)-based minimum variance, provide poor performance when the available snapshots are limited. We propose a Mamba-based beamformer (MambaBF) that leverages an unsupervised deep learning (DL) approach and can be deployed on the user terminal (UT) antenna array, for assisting downlink beamforming and CFI mitigation using only a limited number of available array snapshots as input, and without CSI knowledge. Simulation results demonstrate that MambaBF consistently outperforms conventional beamforming techniques in mitigating interference and maximizing the signal-to-interference-plus-noise ratio (SINR), particularly under challenging conditions characterized by low SINR, limited snapshots, and imperfect CSI.
Index Terms:
Non-Geostationary orbits (NGSOs), co-frequency interference (CFI), Deep Learning (DL), Receive Beamforming
@@ I INTRODUCTION
Satellite communications (SatCom) will play a vital role in next-generation wireless networks by providing service to vast areas that lack terrestrial network coverage, especially with the rapidly growing Low-Earth orbit (LEO) mega-constellations [1]. These constellations are distributed in the Non-Geostationary Satellites orbits (NGSOs) and must manage co-existence strategies for sharing the limited spectrum [2]. Despite the proactive coordination efforts established by the International Telecommunication Union (ITU) among NGSO operators [3], there remains a risk of unintentional co-frequency interference (CFI), especially in the Ku/Ka bands where tens of thousands of NGSOs operate, potentially causing severe quality of service (QoS) degradation [4]. Thus, advanced interference avoidance and mitigation techniques that ensure both low complexity and efficient mitigation of unintentional interference in NGSO spectrum co-existence scenarios are needed.
User terminals (UTs) planar arrays are becoming popular because they can be used to steer the receiver electronically towards the moving satellite line-of-sight [5]. This work focuses on the adaptive beamforming (ABF) techniques on the UT side, where low complexity and fast adaptation are needed (due to the time-varying nature of the scenario). The goal is to generate antenna weights to accurately steer the main lobe towards the signal of interest (SOI) locations while minimizing the antenna gain towards interference locations.
@@
In this paper, we propose an unsupervised DL-based beamformer approach using Mamba state space models (SSMs) layers [10]. These layers effectively capture both spatial and temporal features from the input data that may be relevant to generating the correct beamforming weight vectors. Mamba-based beamformer (MambaBF) can be deployed on the UT side for more energy-efficient training and inference rather than on-board training [10615457], for assisting downlink CFI mitigation in NGSO communications. MambaBF is trained using the available received array snapshots as input to produce a beamforming weight vector that can maximize SINR by nulling interference directions.
@@ II-A Channel and signal models
$\mathbf{v}(\phi,\theta)$ $=\frac{1}{\sqrt{M}}\left[1,\cdots,e^{j\frac{2\pi}{\lambda}d(m_{x}\sin\phi\cos\theta+m_{y}\sin\theta)},\cdots,\right.$ (1)
$\left.e^{j\frac{2\pi}{\lambda}d((M_{x}-1)\sin\phi\cos\theta+(M_{y}-1)\sin\theta)}\right]^{T}\in\mathbb{C}^{M},$
where $d$ is the inter-element spacing, $0\leq m_{x}<M_{x}$ and $0\leq m_{y}<M_{y}$ are the antenna indices in the $xy$-plane, $\phi$ and $\theta$ are the azimuth and the elevation angles respectively, representing the directions of arrival (DoAs). For notational brevity, we define a DoA as $\varphi\triangleq(\phi,\theta)$.
We are modelling a scenario where the $\mathrm{DS}$ transmits a signal ($s_{x}\in\mathbb{C}$) with $\mathbb{E}\{s_{x}s_{x}^{H}\}=1$ to its UT. Eventually, $K(K<M)$ interfering satellites ($\mathrm{IS}_{k},\quad k\in 1,\cdots,K$) are visible to the UT and unintentionally transmit an interference signal ($s_{i,k}\in\mathbb{C}$) towards it. When placing the UT in an obstacle-free elevated space as recommended by many operators (e.g., STARLINK [5]), the line-of-sight (LoS) component becomes dominant, thus the non-LoS (NLoS) propagation components can be ignored. Assuming $\mathrm{DS}$’s Doppler shift is compensated for at the UT, we define the statistical time-varying channel state information (sCSI) model between the $\mathrm{DS}$ satellite and the UT at time instant $t$ and frequency $f_{c,d}$ as follows[12],
@@
where $\eta^{ut}$, $D^{ut}$ are the UPA antenna efficiency and the directivity towards $\varphi$ direction. Note that the channel model adopted in (2) and (3) is applicable over a specific coherence time ”$t$” where the relative positions of the satellites and UT do not change significantly, and the physical channel parameters remain invariant. In this work, we simulate a constitutive sequence of coherence times for the sake of adaptability. Assuming that the UT array at the time instant $t$ can capture $L$ total snapshots, the $l^{th}$ snapshot vector can be expressed as,
@@ II-B Traditional beamforming techniques
where $\mathbf{Y}=\left[\mathbf{y}[1],\mathbf{y}[2],\cdots\mathbf{y}[L]\right]\in\mathbb{C}^{M\times L}$. When $\hat{\mathbf{R}}$ is used to formulate (8), the solution is then called sample matrix inversion (SMI)-based minimum variance beamformer and is given by $\mathbf{w}_{\text{SMI}}=\frac{\hat{\mathbf{R}}^{-1}\mathbf{v}_{d}}{\mathbf{v}_{d}^{H}\hat{\mathbf{R}}^{-1}\mathbf{v}_{d}}$. As $L$ increases, $\hat{\mathbf{R}}$ converges to the theoretical covariance matrix $\mathbf{R}=\mathbf{R}_{d}+\mathbf{R}_{i+n}$ and the corresponding $\mathrm{SINR}$ will approximate the optimal value as $L\rightarrow\infty$ However, as the available snapshots size $L$ decreases, the gap between $\hat{\mathbf{R}}$ and $\mathbf{R}$ increases, which dramatically affects the performance [6]. Here $\mathrm{SINR}$ can be measured based on the $L^{th}$ snapshot as a random variable of a discrete random process [see Eq. 7.90 [6]], instead, we assume average $\mathrm{SINR}$ ($\mathrm{ASINR}$) from the available snapshots,
@@ III-A Model Overview
The real and imaginary parts of $\mathbf{Y}$ are first normalized and then processed by a 1D convolution (conv1D) plus max-pooling (MaxPool1D) across the snapshot dimension for dimensionality reduction. Each Conv1D output is Batch-normalized to stabilize training. After these initial layers, the latent input $\mathbf{x}\in\mathbb{R}^{M_{z}\times L}$, $M_{z}<2M$ is passed through the Mamba-based SSM layers (MmLs), followed by fully connected (FC) layers. The final output is a weight vector $\hat{\mathbf{w}}\in\mathbb{R}^{2M}$, reshaped into $\mathbb{C}^{M}$, then normalized to yield
@@
$\mathbf{f}(\ell)\;=\;\mathbf{W}_{\mathrm{out}}\!\Bigl{(}\mathbf{g}(\ell)\,\odot\,\tilde{\mathbf{x}}(\ell)\Bigr{)},$ (18)
where $\tilde{\mathbf{x}}(\ell)$ is a skip-connection transform of $\mathbf{x}(\ell)$, and $\odot$ denotes element-wise multiplication.
III-C Unsupervised Training Objective
The proposed learning procedure offers unsupervised training, and the loss function is customized to minimize $\mathcal{L}_{\mathrm{SINR}}(n)=-\mathrm{SINR}(n)$ in case of a single snapshot or minimize $\mathcal{L}_{\mathrm{ASINR}}(n)=-\frac{1}{L}\sum_{(n,l)}^{(n,L)}\mathrm{SINR}(l,n)$. Given $\mathbf{w}_{\mathrm{MBF}}(\mathbf{Y}(n))$, we measure $\mathcal{L}(n)$ in terms of the received power from the desired direction vs. interference plus noise. We then perform backpropagation through the MambaBF network using $\mathcal{L}(n+1)$ as our new cost.
@@ IV-A Parameters and Data Generation
Table I parameters were used to generate the interference scenarios in MATLAB, in addition to these parameters, the $\mathrm{DS}$ is assumed to have a pair of DOAs $\left[\phi_{d},\theta_{d}\right]=\left[0,0\right]$, while we assume $\mathrm{ISs}$ transmit severe interference (QoS degradation) could be around side lobe levels (SLLs), so for an array operating in scan angles $\phi,\theta\in\left[-90:90\right]$ the critical angles it is assumed to be $\phi_{k},\theta_{k}\in\left[-40:40\right]$. The UT directivity in (4) points at $\mathrm{DS}$ direction with efficiency $\eta^{ut}=0.99$ and $\mathbf{v}_{d}$ initial weights. This setup produces different $\mathrm{SINR}$ values, particularly low $\mathrm{SINR}$ at highly correlated desired and interference channels (in-line DOAs angles), and higher $\mathrm{SINR}$ values at less correlated channels or at nulled SLLs locations. We will train our models twice for two scenarios, when we have a single snapshot, and when we have multiple snapshots. The multi-snapshot data is generated at a sampling rate $f_{s}=2\times BW_{d}$ producing $200$ snapshots per training data sample ”$n$”, this is for the SMI benchmark where $\mathbf{\hat{R}}$ should be estimated assuming a full rank to ensure a good performance of the matrix inversion.
@@
$M$ $10\times 10$ elements
$K$ 3 satellites
$r_{d}$ , $r_{k}$ 1000, $\left[500~{}\text{to}~{}600\right]$ Km
${P}_{d}$ , ${P}_{k}$ 45 dBW, $40$ dBW
$f_{c,d}$ , $f_{c,i}$ 11.750 GHz
$BW_{d}$ , $BW_{i}$ 50 MHz
$\mathrm{DS}$ $\mathrm{ISs}$ Modcods QPSK, 8QAM
$T^{ut}$ 230 K
@@ IV-C Model Performance
We evaluate the performance of the proposed MambaBF beamformer using off-training (test) samples. Earlier, we assumed UPA has knowledge of the $\mathrm{DS}$’s DOA, thus the initial beamformer $\mathbf{w}_{\text{in}}$ is set to $\mathbf{v}_{d}(\varphi_{d})$. Our goal is to compare various beamformers’ performances by replacing $\mathbf{w}_{\text{in}}$ with a new updated beamformer $\mathbf{w}_{\text{out}}$ and measuring the resulting $\mathrm{SINR}_{\text{out}}(\mathbf{w}_{\text{out}})$ where $\mathbf{w}_{\text{out}}\in[\mathbf{w}_{\mathrm{ZF}},\mathbf{w}_{\mathrm{MRC}}$ $,\mathbf{w}_{\mathrm{SMI}},\mathbf{w}_{\mathrm{MBF}}]$.
Figs. 4 plot $\mathrm{SINR}_{\mathrm{in}}$ versus $\mathrm{SINR}_{\mathrm{out}}$ for both perfect CSI and imperfect CSI conditions. a) Perfect CSI (Left Figure): SMI exhibits poor performance when the number of snapshots (here 200) is limited  [^1] The performance of SMI can improve and get closer to optimal MVDR if the number of snapshots is significantly increased (e.g., 100,000 snapshots), which is substantially higher than the 200 snapshots used for MambaBF.. ZF performs exceptionally well under high-quality CSI by exploiting full knowledge of the desired and interference channels; it fully cancels interference but struggles in negative $\mathrm{SINR}_{\mathrm{in}}$ due to the high correlation between desired and interference channels (i.e., $\mathbf{v}_{d}\approx\mathbf{v}_{k}$). MRC performs the worst among the classical approaches since it maximizes only the desired signal power without explicitly nulling interference. MambaBF consistently outperforms MRC and sometimes outperforms ZF at low $\mathrm{SINR}_{\mathrm{in}}$, converging at higher $\mathrm{SINR}_{\mathrm{in}}$ with minor fluctuations. b) Imperfect CSI (Right Figure): With channel estimation errors, SMI, ZF, and MRC see noticeable performance drops, illustrating their sensitivity to imperfect CSI. The MambaBF model, in contrast, outperforms these methods across both low and high $\mathrm{SINR}_{\mathrm{in}}$, showing robustness under realistic imperfect CSI conditions.
@@ V Conclusion and Future Work
We have presented MambaBF, an unsupervised DL beamforming method that maximizes the SINR and suppresses interference using only the available array snapshots, even with a single snapshot. Numerical results confirm that MambaBF is the most flexible approach, especially under imperfect CSI, where classical methods deteriorate. This highlights MambaBF’s potential for real-world deployments with inevitable channel imperfections. Potential future directions include evaluating MambaBF under more complex scenarios with moving users, as well as exploring larger-scale datasets for improving the beamforming capability to ZF (optimal) level performance.

### v2

@@ Abstract
In this paper, we investigate downlink co-frequency interference (CFI) mitigation in non-geostationary satellite orbits (NGSOs) co-existing systems. Traditional mitigation techniques, such as Zero-forcing (ZF), produce a null towards the direction of arrivals (DOAs) of the interfering signals, but they suffer from high computational complexity due to matrix inversions and required knowledge of the channel state information (CSI). Furthermore, adaptive beamformers, such as sample matrix inversion (SMI)-based minimum variance, provide poor performance when the available snapshots are limited. We propose a Mamba-based beamformer (MambaBF) that leverages an self-supervised deep learning (DL) approach and can be deployed on the user terminal (UT) antenna array, for assisting downlink beamforming and CFI mitigation using only a limited number of available array snapshots as input, and without CSI knowledge. Simulation results demonstrate that MambaBF consistently outperforms conventional beamforming techniques in mitigating interference and maximizing the signal-to-interference-plus-noise ratio (SINR), particularly under challenging conditions characterized by low SINR, limited snapshots, and imperfect CSI.
@@ I INTRODUCTION
Satellite communications (SatCom) will play a vital role in next-generation wireless networks by providing service to vast areas that lack terrestrial network coverage, especially with the rapidly growing Low-Earth orbit (LEO) mega-constellations [1]. These constellations are distributed in the Non-Geostationary Satellite orbits (NGSOs) and must manage co-existence strategies for sharing the limited spectrum [2]. Despite the proactive coordination efforts established by the International Telecommunication Union (ITU) among NGSO operators [3], there remains a risk of unintentional co-frequency interference (CFI), especially in the Ku/Ka bands where tens of thousands of NGSOs operate, potentially causing severe quality of service (QoS) degradation [4]. Thus, advanced interference avoidance and mitigation techniques that ensure both low complexity and efficient mitigation of unintentional interference in NGSO spectrum co-existence scenarios are needed.
Planar arrays are becoming popular user terminals (UT) equipments, because they can be used to steer the receiver electronically towards the moving satellite line of sight [5]. This work focuses on the adaptive beamforming (ABF) techniques on the UT side, where low complexity and fast adaptation are needed (due to the time-varying nature of the scenario). The goal is to generate antenna weights to accurately steer the main lobe towards the signal of interest (SOI) locations while minimizing the antenna gain towards interference locations.
@@
In this paper, we propose an self-supervised DL-based beamformer approach using Mamba state space model (SSM) layers [10]. These layers effectively capture both spatial and temporal features from the input data that may be relevant to generating the correct beamforming weight vectors. Mamba-based beamformer (MambaBF) can be deployed on the UT side for more energy-efficient training and inference rather than on-board training [11], for assisting downlink CFI mitigation in NGSO communications. MambaBF is trained using the available received array snapshots as input to produce a beamforming weight vector that can maximize SINR by nulling interference directions.
@@ II-A Channel and signal models
$\mathbf{v}(\phi,\theta)$ $=\frac{1}{\sqrt{M}}\left[1,\cdots,e^{j\frac{2\pi}{\lambda}a(m_{x}\sin\phi\cos\theta+m_{y}\sin\theta)},\cdots,\right$. (1)
$\left.e^{j\frac{2\pi}{\lambda}a((M_{x}-1)\sin\phi\cos\theta+(M_{y}-1)\sin\theta)}\right]^{T}\in\mathbb{C}^{M},$
where $a$ is the inter-element spacing, $0\leq m_{x}<M_{x}$ and $0\leq m_{y}<M_{y}$ are the antenna indices in the $xy$-plane, $\phi$ and $\theta$ are the azimuth and the elevation angles respectively, representing the directions of arrival (DoAs). For notational brevity, we define a DoA as $\varphi\triangleq(\phi,\theta)$.
We are modelling a scenario where the $\mathrm{DS}$ transmits a signal ($s_{x}\in\mathbb{C}$) with $\mathbb{E}\{s_{x}s_{x}^{H}\}=1$ to its UT. Eventually, $K(K<M)$ interfering satellites ($\mathrm{IS}_{k},\quad k\in 1,\cdots,K$) are visible to the UT and unintentionally transmit an interference signal ($s_{i,k}\in\mathbb{C}$) towards it (the underscore ”ii” and ”dd” will be used to indicate interference satellite and desired satellite parameters respectively). When placing the UT in an obstacle-free elevated space as recommended by many operators (e.g., STARLINK [5]), the line-of-sight (LoS) component becomes dominant, thus the non-LoS (NLoS) propagation components can be ignored. Assuming $\mathrm{DS}$’s Doppler shift is compensated for at the UT, we define the statistical time-varying channel state information (sCSI) model between the $\mathrm{DS}$ satellite and the UT at time instant $t$ and frequency $f_{c,d}$ as follows[13],
@@
where $\eta^{ut}$, $D^{ut}$ are the UPA antenna efficiency and the directivity towards $\varphi$ direction. Note that the channel model adopted in (2) and (3) is applicable over a specific coherence time ”tt” where the relative positions of the satellites and UT do not change significantly, and the physical channel parameters remain invariant. In this work, we simulate a constitutive sequence of coherence times for the sake of adaptability. Assuming that the UT array at the time instant $t$ can capture $L$ total snapshots, the $l^{th}$ snapshot vector can be expressed as,
@@ II-B Traditional beamforming techniques
where $\mathbf{Y}=\left[\mathbf{y}[1],\mathbf{y}[2],\cdots\mathbf{y}[L]\right]\in\mathbb{C}^{M\times L}$. When $\hat{\mathbf{R}}$ is used to formulate (8), the solution is then called sample matrix inversion (SMI)-based minimum variance beamformer and is given by $\mathbf{w}_{\text{SMI}}=\frac{\hat{\mathbf{R}}^{-1}\mathbf{v}_{d}}{\mathbf{v}_{d}^{H}\hat{\mathbf{R}}^{-1}\mathbf{v}_{d}}$. As $L$ increases, $\hat{\mathbf{R}}$ converges to the theoretical covariance matrix $\mathbf{R}=\mathbf{R}_{d}+\mathbf{R}_{i+n}$ and the corresponding $\mathrm{SINR}$ will approximate the optimal value as $L\rightarrow\infty$ . However, as the available snapshots size $L$ decreases, the gap between $\hat{\mathbf{R}}$ and $\mathbf{R}$ increases, which dramatically affects the performance [6]. Here $\mathrm{SINR}$ can be measured based on the $L^{th}$ snapshot as a random variable of a discrete random process [see Eq. 7.90 [6]], instead, we assume average $\mathrm{SINR}$ ($\mathrm{ASINR}$) from the available snapshots,
@@ III-A Model Overview
The real and imaginary parts of $\mathbf{Y}$ are first normalized and then processed by a 1D convolution (conv1D) plus max-pooling (MaxPool1D) across the snapshot dimension for dimensionality reduction. Each Conv1D output is Batch-normalized to stabilize training. After these initial layers, the latent input $\mathbf{x}\in\mathbb{R}^{M_{z}\times L}$, $M_{z}<2M$, where $M_{z}$ is the latent space size, is passed through the Mamba-based SSM layers (MmLs), followed by fully connected (FC) layers. The final output is a weight vector $\hat{\mathbf{w}}\in\mathbb{R}^{2M}$, reshaped into $\mathbb{C}^{M}$, then normalized to yield
@@
[…]
III-C self-supervised Training Objective
The proposed learning procedure offers self-supervised training, and the loss function is customized to minimize $\mathcal{L}_{\mathrm{ASINR}}(n)=-\frac{1}{L}\sum_{(n,l)}^{(n,L)}\mathrm{SINR}(l,n)$, where $n$ is a training sample. Given $\mathbf{w}_{\mathrm{MBF}}(\mathbf{Y}(n))$, we measure $\mathcal{L}(n)$ in terms of the received power from the desired direction vs. interference plus noise. We then perform backpropagation through the MambaBF network using $\mathcal{L}(n+1)$ as our new cost.
@@ IV-A Parameters and Data Generation
Table I parameters were used to generate interference scenarios in MATLAB, in addition to these parameters, for simplification, the UPA is tilted to capture the $\mathrm{DS}$ signal from DOAs $\left[\phi_{d},\theta_{d}\right]=\left[0,0\right]$, while we assume that $\mathrm{ISs}$ transmit severe interference (QoS degradation) could be around side lobe levels (SLLs), so for an array operating in scan angles $\phi,\theta\in\left[-90:90\right]$ the critical angles are assumed to be $\phi_{k},\theta_{k}\in\left[-40:40\right]$. The UT directivity in (4) points at $\mathrm{DS}$ direction with efficiency $\eta^{ut}=0.99$ and $\mathbf{v}_{d}$ initial weights. This setup produces different $\mathrm{SINR}$ values, particularly low $\mathrm{SINR}$ at highly correlated desired and interference channels (in-line DOAs angles), and higher $\mathrm{SINR}$ values at less correlated channels or at nulled SLLs locations. The multi-snapshot data is generated at a sampling rate $f_{s}=2\times$ Signal Bandwidth, producing $200$ snapshots per training data sample ”nn”, this is for the SMI benchmark where $\mathbf{\hat{R}}$ should be estimated assuming a full rank to ensure a good performance of the matrix inversion.
@@
Number of elements ($M$) $10\times 10$ elements
Number of interfering satellites ($K$) 3 satellites
Satellites altitude ($r_{d}$ , $r_{k}$) 1000, $\left[500~\text{to}~600\right]$ Km
Satellites EIRP (${P}_{d}$ , ${P}_{k}$) 45 dBW, $40$ dBW
Signal Bandwidth ($s_{x}$, $s_{i,k}$) 50 MHz
Signal Downlink Frequency ($f_{c,d}$ , $f_{c,i}$) 11.750 GHz
Signal Modcods ($\mathrm{DS}$ & $\mathrm{ISs}$) QPSK, 8QAM
Number of snapshots ($L$) 200200 snapshots
UT equivalent noise temperature ($T^{ut}$) 230 K
@@ IV-C Model Performance
We evaluate the performance of the proposed MambaBF beamformer using off-training (test) samples. Earlier, we assumed UPA has knowledge of the $\mathrm{DS}$’s DOA, thus the initial beamformer $\mathbf{w}_{\text{in}}$ is set to $\mathbf{v}_{d}(\varphi_{d})$, thus the measured SINR using the initial beamformer will be indicated by $\mathrm{SINR}_{\text{in}}$ in this paper. Our goal is to compare various beamformers’ performances by replacing $\mathbf{w}_{\text{in}}$ with a new updated beamformer $\mathbf{w}_{\text{out}}$ and measuring the resulting $\mathrm{SINR}_{\text{out}}(\mathbf{w}_{\text{out}})$ where $\mathbf{w}_{\text{out}}\in[\mathbf{w}_{\mathrm{ZF}},\mathbf{w}_{\mathrm{MRC}}$ $\mathbf{w}_{\mathrm{SMI}},\mathbf{w}_{\mathrm{MBF}}]$.
Figs. 4 plot $\mathrm{SINR}_{\mathrm{in}}$ versus $\mathrm{SINR}_{\mathrm{out}}$ for both perfect CSI and imperfect CSI conditions. a) Perfect CSI (Left Figure): SMI exhibits poor performance when the number of snapshots (here 200) is limited  [^1] The performance of SMI can improve and get closer to optimal MVDR if the number of snapshots is significantly increased (e.g., 100,000 snapshots), which is substantially higher than the 200 snapshots used for MambaBF.. ZF performs exceptionally well under high-quality CSI by exploiting full knowledge of the desired and interference channels; it fully cancels interference but struggles in negative $\mathrm{SINR}_{\mathrm{in}}$ due to the high correlation between desired and interference channels (i.e., $\mathbf{v}_{d}\approx\mathbf{v}_{k}$). MRC performs the worst among the classical approaches since it maximizes only the desired signal power without explicitly nulling interference. MambaBF mostly outperforms MRC and sometimes outperforms ZF at low $\mathrm{SINR}_{\mathrm{in}}$, converging at higher $\mathrm{SINR}_{\mathrm{in}}$ with minor fluctuations. b) Imperfect CSI (Right Figure): With channel estimation errors, SMI, ZF, and MRC see noticeable performance drops, illustrating their sensitivity to imperfect CSI. The MambaBF model, in contrast, outperforms these methods across both low and high $\mathrm{SINR}_{\mathrm{in}}$, showing robustness under realistic imperfect CSI conditions.
@@ V CONCLUSION AND FUTURE WORK
We have presented MambaBF, an self-supervised DL beamforming method that maximizes the SINR and suppresses interference using only the available array snapshots. Numerical results confirm that MambaBF is the most flexible approach, especially under imperfect CSI, where classical methods deteriorate. This highlights MambaBF’s potential for real-world deployments with inevitable channel imperfections. Potential future directions include evaluating MambaBF under more complex scenarios with moving users, as well as exploring larger-scale datasets for improving the beamforming capability to ZF (optimal) level performance.